OpenStudy (anonymous):

Solve: (x - 8)(x + 11) = 0

OpenStudy (campbell_st):

2 solutions when is x - 8 = 0 and x + 11 = 0 they will give the solutions as multiplying by zero is zero

OpenStudy (callisto):

To solve such equation, you need to know that either (x - 8)=0 or (x + 11) = 0 So, (x - 8)=0 or (x + 11) = 0 x-8+8 =0+8 or x+11-11 =0-11 Can you solve them from here?